Grade 11Mechanics

a body covers one third part of its journey with a velocity of 2 metre per second next one third part with a velocity of 3 metre per second and the rest of the journey with a velocity of 6 metre per second then what is the average velocity of the body

Virati
8 Years agoGrade 11

1 Answer

Arun
8 Years ago
Dear student
 
Averagw velocity = (x/3 + x/3 + x/3)/[x/6 + x/9 + x/18]
 
 = x/[3x+ 2x +x/18] = 18 x/6x = 3 m/sec
 
Hope it helps